# A method to benchmark high-dimensional process drift detection

### v1

@@ 5.4. Results
The result of our benchmark study is shown in Figure 7. We again want to mention that our benchmark study is not to show whether one model is superior over another. Instead, its purpose is to highlight how well the TAUC and our synthetization system allows to explore their performance. Surprisingly, the RandomGuessDetector reached the highest AUC score on dataset-3, where it ranges on all three datasets among the last ranks in the TAUC score. The respective predictions of the best detectors are plotted in Figure 9. Moreover, cluster based systems reach good AUC scores, but are not competitive when using the TAUC as quantification measurement (see also Figure 8).
Fig. 7. Benchmark results on dataset-1, dataset-2, and dataset-3
Unsurprisingly, the performance of autoencoders-based systems depends heavily on their hyperparameters, leading to large variations in the AUC score.
Fig. 8. Performance of cluster based systems on dataset-3, ranging among the models with highest AUC, but lowest TAUC.
Fig. 9. Comparison of the prediction $s\in\mathbb{R}^{T}$ of the best detectors on dataset-3.
@@ Acknowledgements
This work is supported by the Hightech Agenda Bavaria. The authors are grateful to Matthias Burkhardt, Fabian Hueber, Kai Müller, and Ulrich Göhner for helpful discussions.
@@
Appendix A Data generation with polynomials
@@
with $w\in\mathbb{R}^{6}.$ We simulate $2000$ process executions and thus sample $2000$ process curves. The shape of each curve is defined by its support points. We are only interested in its curvature in $I=[0,4]$. First, we want to add a condition onto the begin and end of the interval, namely that $f(w,0)=4$ and $f(w,4)=5$. Moreover, we would like to have a global maximum at $x=2$, which means the first order derivative
@@ B.1. Lagged prediction
B.2. Change point detection
@@
B.3. Varying length and position of predicted segments
A situation where the sTAUC coincides with the AUC mostly is in when only one true and predicted drift segment exist (see Figure 14). In cases where the center of the predicted segment coincides with the center of the true segment, the AUC and sTAUC match almost exactly when the length of the predicted segment is varied (see left graphic in Figure 15). If the predicted segment has fixed length that equals the length of the true segment and the position of its center is varied from $50$ to $350$, AUC and sTAUC coincide mostly, but the sTAUC shows a faster rise when the predicted segment overlaps with the true segment due to the effects explained in Section B.1.
@@ B.3. Varying length and position of predicted segments
Appendix C TAUC for trivial detector

### v2

@@ 5.4. Results
The result of our benchmark study is shown in Figure 10. Generally, there is a discrepancy in detectors of the highest AUC and the highest TAUC. More concrete, the larger the number of true drift segments in a dataset is, the larger the discrepancy (see also Figure 11). For instance, the RandomGuessDetector reached the highest AUC score on dataset-1, where it ranges on all three datasets among the last ranks in the TAUC score. On all datasets, autoencoder-based systems reach among the best detectors for both, TAUC and AUC. Those using a multivariate test in their latent space reach better scores than these using an aggregation of multiple uni-variate tests. Although some cluster-based systems archive good AUC scores on dataset dataset-3, none of the benchmarked algorithms is capable to compute a score that can be used to recover the true drift segments, resulting in small TAUC scores for all algorithms (see also Figure 14). The respective predictions over the temporal dimension of the best detectors are shown in Appendix A in more detail, where it also becomes visible that detectors with higher TAUC better recover the true drift segments.
Fig. 10. Benchmark results on dataset-1, dataset-2, and dataset-3.
Fig. 11. Correlation between TAUC and AUC for different number of true drift segments $k$.
6. Conclusion
This work shows how algorithms designed to detect process drifts can be benchmarked in robust and reliable way. We have introduced a scalable and controllable data generation method that creates process curves datasets with drifts and a verified ground truth. In our approach, process curve datasets can be solely generated by modelling the behavior of support points over the temporal axis and using non-linear optimization. We then introduce and study the novel TAUC score which is particularly designed to evaluate the performance of drift detectors on their temporal consistency over sequential process executions. We proved the effectiveness of our approach in a small benchmark study. Our results reveal that existing algorithms often struggle with datasets containing multiple drift segments, underscoring the need for further research.
@@ Acknowledgements
This work is supported by the Hightech Agenda Bavaria. The authors are grateful to Matthias Burkhardt, Fabian Hueber, Kai Müller, and Ulrich Göhner for helpful discussions. We also thank the anonymous referees for helpful comments and suggestions.
@@
Appendix A Predictions of detectors of benchmark study
Figure 12, Figure 13, and Figure 14 show the prediction $s\in\mathbb{R}^{T}$ of the detectors reaching highest TAUC and AUC scores on the individual datasets are shown respectively.
Fig. 12. Best detectors on dataset-1.
Fig. 13. Best detectors on dataset-2.
Fig. 14. Best detectors on dataset-3.
Appendix B Data generation with polynomials
@@
with $w\in\mathbb{R}^{6}.$ We simulate $2000$ process executions and thus sample $2000$ process curves. The shape of each curve is defined by its support points. We are only interested in its curvature in $I=[0,4]$. First, we want to add a condition onto the start and end of the interval, namely that $f(w,0)=4$ and $f(w,4)=5$. Moreover, we would like to have a global maximum at $x=2$, which means the first order derivative
@@ C.1. Lagged prediction
C.2. Change point detection
@@
C.3. Varying length and position of predicted segments
A situation where the sTAUC coincides with the AUC mostly is in when only one true and predicted drift segment exist (see Figure 19). In cases where the center of the predicted segment coincides with the center of the true segment, the AUC and sTAUC match almost exactly when the length of the predicted segment is varied (see left graphic in Figure 20). If the predicted segment has fixed length that equals the length of the true segment and the position of its center is varied from $50$ to $350$, AUC and sTAUC coincide mostly, but the sTAUC shows a faster rise when the predicted segment overlaps with the true segment due to the effects explained in Section C.1.
@@ C.3. Varying length and position of predicted segments
Appendix D TAUC for trivial detector